Circular No. 5812

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993R IN NGC 7742 R. R. Treffers, B. Leibundgut, and A. V. Filippenko, University of

California at Berkeley; and M. W. Richmond, Princeton University, report the discovery of a probable supernova in NGC 7742 (α = 23^{h}41^{m}43^{s}, δ = +10^{o}29′.3, equinox 1950.0). The object was found during the Leuschner Observatory Supernova Search (LOSS), which uses an automated 0.76-m telescope equipped with the Lawrence Berkeley Laboratory CCD camera. SN 1993R was detected in an image obtained on June 2 UT and confirmed on June 3. Inspection of earlier LOSS observations of NGC 7742 reveals the presence of this object since Feb. 15, before it disappeared behind the sun. It was not present in an image obtained on Feb. 2. The supernova is located ∼ 8′′ east and 5′′ north of the galaxy nucleus and is estimated to be of magnitude R ∼ 17.

SUPERNOVAE 1993S AND 1993T IN ANONYMOUS GALAXIES J. Maza, University of Chile; and M. Hamuy, Cerro Tololo Interamer-

ican Observatory, report the discovery of two apparent supernovae by R. Antezana (University of Chile) on 20-min unfiltered IIa-O plates taken by G. Valladares with the CTIO Curtis Schmidt telescope. SN 1993S (m_{pg} ∼ 17.5 on May 26.36 UT) is located 7′′.5 west and 6′′.5 north of the nucleus of a spiral galaxy at α = 22^{h}49^{m}33^{s}.4, δ = −40^{o}34′40′′ (equinox 1950.0). Confirmation of SN 1993S was made by R. Aviles on CCD B, V, and I images obtained on June 5.36 with the CTIO 0.9-m telescope. SN 1993T (m_{pg} ∼ 18 on May 26.38 UT) is located ∼ 18′′.5 west and 13′′.5 north of the nucleus of a galaxy at α = 23^{h}08^{m}07^{s}.7, δ = −45^{o}15′21′′ (equinox 1950.0). Confirmation of SN 1993T was also made by Aviles from a CCD V image obtained on June 5.38.

SUPERNOVA 1993J IN NGC 3031 Further photoelectric photometry by H. Mikuž, B. Dintinjana, and T.

Zwitter, Ljubljana, Slovenia (cf. IAUC 5796): May 31.93 UT, B = 13.45 ± 0.03, V = 12.53 ± 0.01, R = 11.96 ± 0.02, I = 11.65 ± 0.02; June 1.88, 14.13 ± 0.02, 12.62 ± 0.02, 12.03 ± 0.02, 11.75 ± 0.02; 5.90, 13.69 ± 0.06, 12.67 ± 0.01, 12.10 ± 0.04, 11.70 ± 0.04.

1993 June 10 Daniel W. E. Green